OpenStudy (anonymous):

@ganeshie8 @Kristen17 Solve for x. Enter your answer below as an improper fraction in lowest terms, using the slash ( / ) as the fraction bar. x-6(9/10)=5/12

OpenStudy (charlotte123):

x - 27/5 = 5/12 <---- + 27/5 + 27/5 Thus x = 349/60.

OpenStudy (anonymous):

x - 6(9/10) = 5/12 x - 27/5 = 5/12 x = 349/60

OpenStudy (charlotte123):

(5 x)/5-(27)/5 = 5/12 (5 x)/5-(27)/5 = (5 x-27)/5: (5 x-27)/5 = 5/12 (5 (5 x-27))/5 = 5×5/12 5×5/12 = (5×5)/12: (5 (5 x-27))/5 = (5×5)/12 (5 (5 x-27))/5 = 5/5×(5 x-27) = 5 x - 27 5 x - 27 = (5×5)/12 5×5 = 25: 5 x-27 = 25/12

OpenStudy (charlotte123):

5 x + (27-27) = 25/12+27 27 - 27 = 0 5 x = 25/12+27 5 x = 25/12+(12×27)/12 12×27 = 324 5 x = 25/12 + 324/12 25/12 + 324/12 = (25+324)/12: 5 x = (25+324)/12 25+324 = 349 5 x = 349/12

OpenStudy (charlotte123):

x = 349/(12×5) 12×5 = 60

OpenStudy (anonymous):

What in the world did you just do????????????????

OpenStudy (charlotte123):

Math :)

OpenStudy (anonymous):

Yea you did, my way is much easier

OpenStudy (charlotte123):

LOL

OpenStudy (anonymous):

@mayafrancis are there medals involved??

OpenStudy (anonymous):

Ty

OpenStudy (charlotte123):

No problem! Thank you as well

OpenStudy (anonymous):

Np ;)

OpenStudy (charlotte123):

LOLLLLLLLLLLLLLLLL

undeadknight26 (undeadknight26):

ma head hurts...

OpenStudy (anonymous):

Your welcome @mayafrancis

OpenStudy (anonymous):

Lol @undeadknight26 were you reading @charlotte123 response

OpenStudy (charlotte123):

THE POWERS OF MATH NO REGRETS

OpenStudy (anonymous):

XD

undeadknight26 (undeadknight26):

yes yes i was...

OpenStudy (anonymous):

LOL!!!! XD Ikr